OpenStudy (anonymous):

Solve for L. L=110+25 cos(4pi)

OpenStudy (campbell_st):

do you know the value of \[\cos(4\pi)\]

OpenStudy (unklerhaukus):

it might help to know that cosine repeats its values every 2π radians (360° degrees) so \(\cos(\theta)=\cos(\theta+2\pi)=\cos(\theta+4\pi)=\dots=\cos(\theta+2n\pi)\) where \(n\) is an integer: \(n\in\mathbb Z\)

OpenStudy (unklerhaukus):

so,\[\cos(4\pi)=\cos(2\pi)=\cos(0\pi)=\cos(0)\] do you know the value of the cosine of zero?